We are working with domains, ranges and functions.
Here is one problem I've already done:
Let f(x)=x^2-1 and g(x)= 1-2x . Find the indicated values.
g(g(2))
g(2)=1-2(2)= -3
g(3)=1-2(-3)= 7
g= -3 and 7

I'm at a lose for the fraction ones, like:

Let f(x)=x^2-1 and g(x)= 1-2x. Find the indicated values.

f(1/g(3))

Answer by jsmallt9(3296)

I'm at a loss for the fraction ones, like:

Let f(x)=x^2-1 and g(x)= 1-2x. Find the indicated values.

f(1/g(3))
**************
g(3) = 1 - 2(3)
g(3) = 1 - 6
g(3) = -5

1/g(3) = 1/-5

f(1/g(3) = f(1/-5) = (1/-5)^2 - 1 = 1/25 - 1 = 1/25 - 25/25 = -24/25